Using the world population formula P = 6.9(1.011)t, where t is the number of



years after 2010 and P is the world population in billions of people, estimate:



a) the population in the year 2030 to the nearest hundred million, and



b) by what year will the population be double from 2010?


1
Expert's answer
2021-12-09T03:23:43-0500
"P = 6.9(1.011)^t"

a)


"t=2030-2010=20(years)"

"P = 6.9(1.011)^{20}=8.6"

The world population will be 8.6 billions of people in 2030.


b)


"P(t_1) = 6.9(1.011)^{t_1}=2(6.9)"

"(1.011)^{t_1}=2"

"t_1=\dfrac{\ln2}{\ln1.011}"

"t_1=63.36\ years"

"2010+64=2074"

The population will be double by 2074.